# Revisiting Isocurvature Bounds on the Minimal QCD Axion

### v1

@@ 3 Enhanced $f_{a}$ during inflation and parametric resonance
$\Phi(t,\mathbf{x})=\frac{1}{\sqrt{2}}\phi(t,\mathbf{x})\exp\left(i\,\frac{a(t,\mathbf{x})}{f_{a}}\right)\,.$ (3.1)
[…]
$\mathscr{L}_{\text{minimal}}[\Phi]=|\partial_{\mu}\Phi|^{2}-\big{(}\lambda_{\Phi}|\Phi|^{4}-\lambda_{\Phi}f_{a}^{2}|\Phi|^{2}\big{)}+2\xi R|\Phi|^{2}+N_{\textsc{dw}}\frac{a}{f_{a}}\frac{\alpha_{\text{s}}}{8\pi}G_{A\mu\nu}\widetilde{G}^{\mu\nu}_{A},$ (3.2)
with the last term solving the strong CP problem. Note that we have included the $R|\Phi|^{2}$ term because it is allowed by all symmetries, as indeed it must be since $R$ and $|\Phi|^{2}$ are separately allowed (as they must be or we would not have gravity or break PQ). Choosing not to write this term is equivalent to an infinite fine-tuning of its coefficient $\xi$ to zero. In this paper, we focus on the natural possibility that $\xi>0$, so that the potential $V(\Phi)\supset-2\xi R|\Phi^{2}|$ features a negative quadratic term when $R>0$ [45, 46]. During inflation this term takes the value
[…]
so indeed it gives a negative contribution to the mass-squared of $\Phi$. As we discuss in Appendix B, other couplings are also well-motivated, such as couplings to the inflaton field or its derivative [16] or other fields [22, 26, 24, 68, 69], and would lead to qualitatively similar dynamics to what we discuss in this section. [^7] Different approaches to solve the axion isocurvature problem involve the introduction of an axion mass during inflation from PQ-breaking terms [47, 43, 70, 71], or the Witten effect induced by magnetic monopoles [23, 7]. We will stick to the Jordan frame for simplicity for the remainder of this paper, as the physical description is independent of the chosen frame (see e.g. [72, 73]).
@@
$\lambda_{\Phi}\left(f_{a}^{\text{(inf)}}\right)^{2}\equiv\lambda_{\Phi}f_{a}^{2}+2\xi R\quad\longrightarrow\quad f_{a}^{\text{(inf)}}=\sqrt{f_{a}^{2}+\frac{2\xi R}{\lambda_{\Phi}}}\ \overset{f_{a}^{\text{(inf)}}\gg f_{a}}{\longrightarrow}\ \sqrt{\frac{24\xi}{\lambda_{\Phi}}}\,H_{I}$ (3.4)
@@
During inflation, the saxion minimum would then lie at $\langle\phi\rangle=f_{a}^{\text{(inf)}}/\sqrt{2}$, which for $\xi>0$ lies at a greater distance than $f_{a}$ from the origin. Towards the end of inflation, $R$ decreases (both due to the decrease of $H$, and to the increasing energy fraction in relativistic species whose contribution to $R$ is loop-suppressed), and the saxion field value returns back to $f_{a}$. Having a larger effective $f_{a}^{\text{(inf)}}$ during inflation naturally suppresses isocurvature perturbations as in Eq. 2.6.
@@
The idea that an enhancement of $f_{a}$ during inflation to a larger $f_{a}^{\text{(inf)}}$ might reduce the spread $\sigma_{\theta}$ of the axion fluctuations in Eq. 2.3 and alleviate the isocurvature bound, dates back to [44], and was revived in [45, 46, 47]. This possibility was reassessed in particular after the claim of detection of primordial B-modes in the CMB in 2014, which would have excluded the pre-inflationary option in the minimal scenario using the claimed isocurvature bounds (see Fig. 2).
@@
The mechanism of an enhanced $f_{a}^{\text{(inf)}}$ to suppress isocurvature can be jeopardised by the dynamics of $\Phi$ after the end of inflation. As the saxion field rolls towards the new minimum at $f_{a}\ll f_{a}^{\text{(inf)}}$, the evolution of $\phi$ might be not adiabatic, and the Hubble friction might be not sufficient to dissipate the kinetic energy of the saxion when it reaches the potential minimum. It can then oscillate back and forth in the wine-bottle potential, possibly well above the origin. The non-linear coupling to the axion can then give rise to a phenomenon of parametric resonance [74, 75, 76]. The oscillations of the saxion field can act as a driving term for modes of the axion with physical momentum around the Hubble rate close to the end of inflation, and drive their exponential growth. If this phenomenon goes on for a sufficient time, the fluctuations of the angular mode can grow enough to cover the whole compact direction, forming axion strings (potentially with domain walls) and restoring non-thermally the $U(1)$ symmetry: this dynamics was analysed in general terms in [77, 78, 79, 80], and applied to the specific case of $U(1)_{\textsc{pq}}$ after a temporary enhancement of $f_{a}^{\text{(inf)}}$ in [48, 25, 47, 49, 34, 36, 27].
@@
Let us comment on the axion angular speed. In our setup, the complex $\Phi$ field sits in its inflationary vev $f_{a}^{\text{(inf)}}$ throughout the whole inflationary history, so that any initial velocity (either radial or angular) is erased. We also do not introduce any PQ violating operator which might source a velocity $\dot{a}$ for the axion, as it is instead commonly assumed in the literature about the kinetic misalignment mechanism for the axion [42, 35, 81, 38, 82, 37, 83, 41, 40]. As recently discussed in [84], an initial angular speed can contribute to the parametric resonance and induce a non-thermal symmetry restoration for $\Phi$. This is not a concern for our scenario.
@@
We now review the dynamics of saxion and axion in Sections 4 and A, to assess the regions of parameter space (in particular, a small $\lambda_{\Phi}$) where the effect of $\xi$ in Eq. 3.4 suppresses $\sigma_{\theta}$ and avoids isocurvature constraints, without exciting a significant parametric resonance for the axion field.
@@ 4.1 Equations of motion of the system
We take Eq. 3.2 as the Lagrangian density for the complex $\Phi$ field. The standard parametrisation of Eq. 3.1 into radial and angular mode is not ideal to compute the evolution of the angular mode in Fourier space, because of non-linearities that couple the saxion and axion in the kinetic term (which reads in this basis $|\partial_{\mu}\Phi|^{2}=\tfrac{1}{2}(\partial_{\mu}\phi)^{2}+\tfrac{1}{2}\phi^{2}(\partial_{\mu}\theta)^{2}$), so that there is no perturbative limit in which these non-linear terms can be neglected.
Another possibility is to decompose the field in Cartesian coordinates [48], so that non-linear terms only arise from the quartic term of the potential (whose coefficient $\lambda_{\Phi}$ is very small in the parameter space that we consider). In the Cartesian parametrization,
$\Phi(t,\mathbf{x})=\frac{X(t,\mathbf{x})+iY(t,\mathbf{x})}{\sqrt{2}}$ (4.1)
@@
$\ddot{X}+3H\dot{X}-\frac{1}{a^{2}}\nabla^{2}X+\lambda_{\Phi}X^{3}-\lambda_{\Phi}\left({f_{a}^{\text{(inf)}}}^{2}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}-Y^{2}}\right)X$ $=0\,,$ (4.2a)
$\ddot{Y}+3H\dot{Y}-\frac{1}{a^{2}}\nabla^{2}Y{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}+\lambda_{\Phi}Y^{3}}-\lambda_{\Phi}\left({f_{a}^{\text{(inf)}}}^{2}-X^{2}\right)Y$ $=0\,.$ (4.2b)
We mark in grey the terms that are negligible in our scenario when the parametric resonance is not significant, that is $Y\ll X,f_{a}^{\text{(inf)}}$. From Eq. 4.2, we can see that, as long as $f_{a}^{\text{(inf)}}$ evolves slowly, the solution for the radial mode is $X\approx f_{a}^{\text{(inf)}}$, which cancels the non-linear coupling between $X$ and $Y$ in Eq. 4.2b for $Y$.
@@
$X(t,\mathbf{x})=\overline{X}(t)+\int\frac{\mathrm{d}^{3}x}{(2\pi)^{3/2}}x_{k}(t)e^{i\mathbf{k}\cdot\mathbf{x}}\,,\quad Y(t,\mathbf{x})=\overline{Y}(t)+\int\frac{\mathrm{d}^{3}x}{(2\pi)^{3/2}}y_{k}(t)e^{i\mathbf{k}\cdot\mathbf{x}}\,.$ (4.3)
@@
$\ddot{\overline{X}}+3H\dot{\overline{X}}+\lambda_{\Phi}\overline{X}^{3}-\lambda_{\Phi}\left({f_{a}^{\text{(inf)}}}^{2}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}-\overline{Y}^{2}}\right)\overline{X}$ $=0\,,$ (4.4a)
$\ddot{\overline{Y}}+3H\dot{\overline{Y}}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}+\lambda_{\Phi}\overline{Y}^{3}}-\lambda_{\Phi}\left({f_{a}^{\text{(inf)}}}^{2}-\overline{X}^{2}\right)\overline{Y}$ $=0\,.$ (4.4b)
In the limit of very gradual evolution of the Ricci scalar towards the end of inflation, and initial conditions $\overline{X}=f_{a}^{\text{(inf)}}$ and $\dot{\overline{Y}}=0$, the background $\overline{X}$ field just tracks its minimum and Eq. 4.4 simplifies to
$\overline{X}(t)\approx f_{a}^{\text{(inf)}}(t)=\sqrt{f_{a}^{2}+(2\xi/\lambda_{\Phi})R(t)}\,,\quad\overline{Y}\approx 0\,,$ (4.5)
@@
$\lambda_{\Phi}X^{3}$ $=\lambda_{\Phi}\overline{X}^{3}+3\lambda_{\Phi}\overline{X}^{2}\int_{k}x_{k}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}+3\lambda_{\Phi}\overline{X}\int_{k}\int_{q}x_{k}x_{q}+\lambda_{\Phi}\int_{k}\int_{q}\int_{p}x_{k}x_{q}x_{p}}$ (4.6)
@@
$\ddot{x}_{k}+3H\dot{x}_{k}+\left(\frac{k^{2}}{a^{2}}+\lambda_{\Phi}\left(3\overline{X}^{2}-{f_{a}^{\text{(inf)}}}^{2}\right)\right)x_{k}$ $=0\,,$ (4.7a)
$\ddot{y}_{k}+3H\dot{y}_{k}+\left(\frac{k^{2}}{a^{2}}+\lambda_{\Phi}\left(\overline{X}^{2}-{f_{a}^{\text{(inf)}}}^{2}\right)\right)y_{k}$ $=0\,.$ (4.7b)
Eq. 4.7b nicely highlights the term responsible for the growth of perturbations via parametric (or tachyonic) resonance, which is only negligible if the evolution of the radial mode is almost adiabatic, $\overline{X}(t)\approx f_{a}^{\text{(inf)}}(t)$ as in Eq. 4.5. In the regime when the growth of the angular direction (approximately coinciding with $Y$) is under control, then the fluctuations of the radial direction in Eq. 4.7a remain small as well.

### v2

@@ 3 Enhanced $f_{a}$ during inflation and parametric resonance
$\Phi(t,\mathbf{x})=\frac{1}{\sqrt{2}}\phi(t,\mathbf{x})\exp\left(i\,\frac{a(t,\mathbf{x})}{f_{a}}\right)\,.$ (9)
We refer to the radial degree of freedom $\phi$ as the saxion field, while the angular direction $a$ is the axion. The minimal Lagrangian [^6] Our convention for the signature of the metric is $\text{diag}(+---)$. for $\Phi$ that spontaneously breaks $U(1)_{\textsc{pq}}$, is
$\mathscr{L}_{\text{minimal}}[\Phi]=|\partial_{\mu}\Phi|^{2}-\big(\lambda_{\Phi}|\Phi|^{4}-\lambda_{\Phi}f_{a}^{2}|\Phi|^{2}\big)+2\xi R|\Phi|^{2}+N_{\textsc{dw}}\frac{a}{f_{a}}\frac{\alpha_{\text{s}}}{8\pi}G_{A\mu\nu}\widetilde{G}^{\mu\nu}_{A},$ (10)
with the last term solving the strong CP problem. Note that we have included the $R|\Phi|^{2}$ term because it is allowed by all symmetries, as indeed it must be since $R$ and $|\Phi|^{2}$ are separately allowed (as they must be or we would not have gravity or break PQ). Choosing not to write this term is equivalent to an infinite fine-tuning of its coefficient $\xi$ to zero. In this paper, we focus on the natural possibility that $\xi>0$, so that the potential $V(\Phi)\supset-2\xi R|\Phi^{2}|$ features a negative quadratic term when $R>0$ Folkerts:2013tua; Fairbairn:2014zta. During inflation this term takes the value
$\mathscr{L}_{\text{minimal}}[\Phi]\supset 2\xi R|\Phi|^{2}\overset{\text{inflation}}{\longrightarrow}24\xi H_{I}^{2}|\Phi|^{2}\,,\qquad(\xi>0)$ (11)
so indeed it gives a negative contribution to the mass-squared of $\Phi$. As we discuss in Appendix B, other couplings are also well-motivated, such as couplings to the inflaton field or its derivative Bao:2022hsg or other fields Jeong:2013xta; Nakayama:2015pba; Takahashi:2015waa; Ballesteros:2016xej; Ramazanov:2022kbd, and would lead to qualitatively similar dynamics to what we discuss in this section. [^7] Different approaches to solve the axion isocurvature problem involve the introduction of an axion mass during inflation from PQ-breaking terms Higaki:2014ooa; Kearney:2016vqw; Co:2023mhe; Berbig:2024ufe; Carrasco:2025rud, or the Witten effect induced by magnetic monopoles Kawasaki:2015lpf; Nomura:2015xil; Banerjee:2024ykz. We will stick to the Jordan frame for simplicity for the remainder of this paper, as the physical description is independent of the chosen frame (see e.g. Catena:2006bd; Postma:2014vaa). For our purposes, the choice of a specific model of inflation is irrelevant. When we discuss later the Starobinsky model, we assume that potential as a stand-in for a particular cold inflation potential. If one defines it in the Jordan frame, the corresponding conformal transformation might further affect the saxion potential by reducing the value of $f_{a}^{\text{(inf)}}$ with respect to $f_{a}$, as described in Rigouzzo:2025hza.
@@
$\lambda_{\Phi}\left(f_{a}^{\text{(inf)}}\right)^{2}\equiv\lambda_{\Phi}f_{a}^{2}+2\xi R\quad\longrightarrow\quad f_{a}^{\text{(inf)}}=\sqrt{f_{a}^{2}+\frac{2\xi R}{\lambda_{\Phi}}}\ \overset{f_{a}^{\text{(inf)}}\gg f_{a}}{\longrightarrow}\ \sqrt{\frac{24\xi}{\lambda_{\Phi}}}\,H_{I}$ (12)
@@
During inflation, the saxion minimum would then lie at $\langle\phi\rangle=f_{a}^{\text{(inf)}}/\sqrt{2}$, which for $\xi>0$ lies at a greater distance than $f_{a}$ from the origin. Towards the end of inflation, $R$ decreases (both due to the decrease of $H$, and to the increasing energy fraction in relativistic species whose contribution to $R$ is loop-suppressed), and the saxion field value returns back to $f_{a}$. Having a larger effective $f_{a}^{\text{(inf)}}$ during inflation naturally suppresses isocurvature perturbations as in Eq. 6.
@@
The idea that an enhancement of $f_{a}$ during inflation to a larger $f_{a}^{\text{(inf)}}$ might reduce the spread $\sigma_{\theta}$ of the axion fluctuations in Eq. 3 and alleviate the isocurvature bound, dates back to Linde:1991km, and was revived in Kasuya:1996ns; Kawasaki:2008jc; Folkerts:2013tua; Fairbairn:2014zta; Kearney:2016vqw. This possibility was reassessed in particular after the claim of detection of primordial B-modes in the CMB in 2014, which would have excluded the pre-inflationary option in the minimal scenario using the claimed isocurvature bounds (see Fig. 2).
@@
The mechanism of an enhanced $f_{a}^{\text{(inf)}}$ to suppress isocurvature can be jeopardised by the dynamics of $\Phi$ after the end of inflation. As the saxion field rolls towards the new minimum at $f_{a}\ll f_{a}^{\text{(inf)}}$, the evolution of $\phi$ might be not adiabatic, and the Hubble friction might be not sufficient to dissipate the kinetic energy of the saxion when it reaches the potential minimum. It can then oscillate back and forth in the wine-bottle potential, possibly well above the origin. The non-linear coupling to the axion can then give rise to a phenomenon of parametric resonance Shtanov:1994ce; Kofman:1995fi; Kofman:1997yn. The oscillations of the saxion field can act as a driving term for modes of the axion with physical momentum around the Hubble rate close to the end of inflation, and drive their exponential growth. If this phenomenon goes on for a sufficient time, the fluctuations of the angular mode can grow enough to cover the whole compact direction, forming axion strings (potentially with domain walls) and restoring non-thermally the $U(1)$ symmetry: this dynamics was analysed in general terms in Tkachev:1995md; Tkachev:1998dc; Kasuya:1997ha; Kasuya:1999hy, and applied to the specific case of $U(1)_{\textsc{pq}}$ after a temporary enhancement of $f_{a}^{\text{(inf)}}$ in Kawasaki:2013iha; Harigaya:2015hha; Kearney:2016vqw; Kobayashi:2016qld; Ema:2017krp; Co:2017mop; Co:2020dya; Ballesteros:2021bee.
@@
Let us comment on the axion angular speed. In our setup, the complex $\Phi$ field sits in its inflationary vev $f_{a}^{\text{(inf)}}$ throughout the whole inflationary history, so that any initial velocity (either radial or angular) is erased. We also do not introduce any PQ violating operator which might source a velocity $\dot{a}$ for the axion, as it is instead commonly assumed in the literature about the kinetic misalignment mechanism for the axion Co:2019wyp; Co:2019jts; Co:2021rhi; Eroncel:2022vjg; Jaeckel:2016qjp; Domcke:2022wpb; Berges:2019dgr; Eroncel:2025qlk; Bodas:2025eca. As recently discussed in Fedderke:2025sic, an initial angular speed can contribute to the parametric resonance and induce a non-thermal symmetry restoration for $\Phi$. This is not a concern for our scenario.
@@
We now review the dynamics of saxion and axion in Sections 4 and A, to assess the regions of parameter space (in particular, a small $\lambda_{\Phi}$) where the effect of $\xi$ in Eq. 12 suppresses $\sigma_{\theta}$ and avoids isocurvature constraints, without exciting a significant parametric resonance for the axion field.
@@ 4.1 Equations of motion of the system
We take Eq. 10 as the Lagrangian density for the complex $\Phi$ field. The standard parametrisation of Eq. 9 into radial and angular mode is not ideal to compute the evolution of the angular mode in Fourier space, because of non-linearities that couple the saxion and axion in the kinetic term (which reads in this basis $|\partial_{\mu}\Phi|^{2}=\tfrac{1}{2}(\partial_{\mu}\phi)^{2}+\tfrac{1}{2}\phi^{2}(\partial_{\mu}\theta)^{2}$), so that there is no perturbative limit in which these non-linear terms can be neglected.
Another possibility is to decompose the field in Cartesian coordinates Kawasaki:2013iha, so that non-linear terms only arise from the quartic term of the potential (whose coefficient $\lambda_{\Phi}$ is very small in the parameter space that we consider). In the Cartesian parametrization,
$\Phi(t,\mathbf{x})=\frac{X(t,\mathbf{x})+iY(t,\mathbf{x})}{\sqrt{2}}$ (13)
@@
$\ddot{X}+3H\dot{X}-\frac{1}{a^{2}}\nabla^{2}X+\lambda_{\Phi}X^{3}-\lambda_{\Phi}\left({f_{a}^{\text{(inf)}}}^{2}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}-Y^{2}}\right)X$ $=0\,$, (14a)
$\ddot{Y}+3H\dot{Y}-\frac{1}{a^{2}}\nabla^{2}Y{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}+\lambda_{\Phi}Y^{3}}-\lambda_{\Phi}\left({f_{a}^{\text{(inf)}}}^{2}-X^{2}\right)Y$ $=0\,.$ (14b)
We mark in grey the terms that are negligible in our scenario when the parametric resonance is not significant, that is $Y\ll X,f_{a}^{\text{(inf)}}$. From Eq. 14, we can see that, as long as $f_{a}^{\text{(inf)}}$ evolves slowly, the solution for the radial mode is $X\approx f_{a}^{\text{(inf)}}$, which cancels the non-linear coupling between $X$ and $Y$ in Eq. 14b for $Y$.
@@
$X(t,\mathbf{x})=\overline{X}(t)+\int\frac{\mathrm{d}^{3}x}{(2\pi)^{3/2}}x_{k}(t)e^{i\mathbf{k}\cdot\mathbf{x}}\,,\quad Y(t,\mathbf{x})=\overline{Y}(t)+\int\frac{\mathrm{d}^{3}x}{(2\pi)^{3/2}}y_{k}(t)e^{i\mathbf{k}\cdot\mathbf{x}}\,.$ (15)
@@
$\ddot{}\overline{X}+3H\dot{}\overline{X}+\lambda_{\Phi}\overline{X}^{3}-\lambda_{\Phi}\left({f_{a}^{\text{(inf)}}}^{2}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}-\overline{Y}^{2}}\right)\overline{X}$ $=0\,$, (16a)
$\ddot{}\overline{Y}+3H\dot{}\overline{Y}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}+\lambda_{\Phi}\overline{Y}^{3}}-\lambda_{\Phi}\left({f_{a}^{\text{(inf)}}}^{2}-\overline{X}^{2}\right)\overline{Y}$ $=0\,.$ (16b)
In the limit of very gradual evolution of the Ricci scalar towards the end of inflation, and initial conditions $\overline{X}=f_{a}^{\text{(inf)}}$ and $\dot{}\overline{Y}=0$, the background $\overline{X}$ field just tracks its minimum and Eq. 16 simplifies to
$\overline{X}(t)\approx f_{a}^{\text{(inf)}}(t)=\sqrt{f_{a}^{2}+(2\xi/\lambda_{\Phi})R(t)}\,,\quad\overline{Y}\approx 0\,$, (17)
@@
$\lambda_{\Phi}X^{3}$ $=\lambda_{\Phi}\overline{X}^{3}+3\lambda_{\Phi}\overline{X}^{2}\int_{k}x_{k}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}+3\lambda_{\Phi}\overline{X}\int_{k}\int_{q}x_{k}x_{q}+\lambda_{\Phi}\int_{k}\int_{q}\int_{p}x_{k}x_{q}x_{p}}$ (18)
@@
$\ddot{x}_{k}+3H\dot{x}_{k}+\left(\frac{k^{2}}{a^{2}}+\lambda_{\Phi}\left(3\overline{X}^{2}-{f_{a}^{\text{(inf)}}}^{2}\right)\right)x_{k}$ $=0\,$, (19a)
$\ddot{y}_{k}+3H\dot{y}_{k}+\left(\frac{k^{2}}{a^{2}}+\lambda_{\Phi}\left(\overline{X}^{2}-{f_{a}^{\text{(inf)}}}^{2}\right)\right)y_{k}$ $=0\,.$ (19b)
Eq. 19b nicely highlights the term responsible for the growth of perturbations via parametric (or tachyonic) resonance, which is only negligible if the evolution of the radial mode is almost adiabatic, $\overline{X}(t)\approx f_{a}^{\text{(inf)}}(t)$ as in Eq. 17. In the regime when the growth of the angular direction (approximately coinciding with $Y$) is under control, then the fluctuations of the radial direction in Eq. 19a remain small as well.